find the equation for the line with y-intercept 3 that is perpendicular to the line y=2/3X-4.
i know that the slope of the line must be the neg reciprocal of the orginal one.
but after that it gets fuzzy

Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(48558)

find the equation for the line with y-intercept 3 that is perpendicular to the line y=2/3X-4.
Right. The slope you want is -3/2 and you are given the y-intercept as 3
That's all you need to know.
The equation is y=(-3/2)x+3
Cheers,
Stan H.

Answer by tutorcecilia(2144)

You're correct, so the second line will have a slope = -3/2
Since the y-intercept = 3, b=3.
Now, you have enough information to plug into the formula:
y = mx + b
y =(-3/2)x + 3